If 40L of an acid solution is 75% acid, how much pure acid is there in the mixture?
Here is what I have tried? 40(75)+100x=85(40+x)
3000+100x=3400+40x
400=60x
x=6.66666

Answer by ptaylor(2198)

If 40L of an acid solution is 75% acid, how much pure acid is there in the mixture?
Here is what I have tried? 40(75)+100x=85(40+x)
3000+100x=3400+40x
400=60x
x=6.66666
OK. THERE IS 75% PURE ACID AND 25% OTHERSTUFF. SO ALL WE NEED TO DO IS FIGURE OUT WHAT 75% OR THREE-FOURTHS OF 40 IS. WELL, 1/2 OF 40 IS 20; 1/4 OF 40 IS 10 AND 3/4 OR 75% OF 40 IS 30.
(0.75*40)=30L-----------ANS
HOPE THIS HELPS---PTAYLOR